[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Nov 13, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Nov 13 06:07:54 UTC 2017


HSU Shift Log for Mon Nov 13 2017 01:07:48 GMT-0500 (EST)

Shift: A

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: Still unable to enter genesis probs for EPAC systems; no  
issues with entering probs for ATLC systems.

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: None.

Miscellaneous issues: None.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171113/4cd7c5b2/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list